# Collaborative and parametric insurance on the Ethereum blockchain

### v1

@@ 3 A state description of the smart contract
A smart contract defines a system whose state evolves based on the actions of blockchain users. In our parametric insurance example, these users include the contract owner, the surplus providers, and the policyholders (see Section 2.1). Participants interact with the smart contract through functions that modify the state of the system. The remainder of this section describes each possible action and its impact on the stochastic processes introduced in Section 2.2.
3.1 Funding the contract
@@
Assume that an investor transfers $x$ ETH to the contract at time $t>0$. Both the balance of the contract and the surplus increase as follows:
@@
where $h>0$ is an infinitesimal time lapse. If the investor is new, then $M_{t+h}=M_{t}+1$; otherwise, $M_{t+h}=M_{t}$. The $j$-th component ($j\in\{1,\ldots,M_{t+h}\}$) of the vector $(Y_{t})_{t\geq 0}$ is updated as:
@@
3.2 Underwrite a parametric insurance policy
A customer wishes to purchase an insurance contract $(A_{T},l)$ at time $t=S$. The premium for such a parametric insurance contract is given by:
@@ 3.2 Underwrite a parametric insurance policy
where $\theta=\mathbb{P}(A_{T})$. The contract can only be underwritten if the surplus of the smart contract is sufficient to cover the resulting increase in the solvency capital requirement. The liability and the cumulative premium collected would then be updated to:
$\widetilde{L}_{t}=L_{t}+l\cdot\mathbb{I}_{A_{T}},\quad\widetilde{\Pi}_{t+h}=\Pi_{t}+\pi.$
@@
$\widetilde{\text{SCR}}_{t}=\text{Quantile}(\widetilde{L}_{t}-\widetilde{\Pi}_{t};\alpha),$
@@
If condition (3) holds, the number of contracts, liability, and premium collected are updated as follows:
$N_{t+h}=N_{t}+1,\quad L_{t+h}=L_{t}+l_{N_{t+h}}\cdot\mathbb{I}_{A_{T_{N_{t+h}}}},\quad\Pi_{t+h}=\Pi_{t}+(1+\eta_{t})\cdot l_{N_{t+h}}\cdot\theta_{N_{t+h}}.$
@@
$\text{SCR}_{t+h}=\text{Quantile}(L_{t+h}-\Pi_{t+h};\,\alpha_{\text{SCR}}),\quad\text{MCR}_{t+h}=\text{Quantile}(L_{t+h}-\Pi_{t+h};\,\alpha_{\text{MCR}}),$
using the recursive method based on a normal approximation of the distribution of $L_{t}$ described in Section 4.
@@
to the vector $\Xi_{t}$ that gathers the statuses of the insurance policies. The balance of the contract is updated as
$B_{t+h}=B_{t}+(1+\eta_{t})\cdot\theta\cdot l,$
while the surplus $X_{t}$ remains unchanged. The policyholder receives:
$y=\frac{\eta_{t}^{(2)}\cdot\theta\cdot l}{r_{t}},$
in protocol tokens. If the policyholder is neither an investor nor a former policyholder, the number of token holders increases as
$M_{t+h}=M_{t}+1,$
otherwise, $M_{t+h}=M_{t}$. A component $j\in\{1,\ldots,M_{t+h}\}$ of the vector $Y_{t}$ is updated as
$Y_{t+h}^{(j)}=Y_{t}^{(j)}+y.$
Finally, the exchange rate becomes
$r_{t+h}=\frac{X_{t+h}}{\bar{Y}_{t+h}}=\frac{X_{t}}{\bar{Y}_{t+h}+y}.$
3.3 Withdrawing from the smart contract and burning tokens
An investor or policyholder $j\in\{1,\ldots,M_{t}\}$ may withdraw some of their funds by burning protocol tokens at time $t>0$. This action is subject to the conditions that the participant holds enough protocol tokens $Y^{(j)}_{t}$, the smart contract has a sufficient balance $B_{t}$, and the surplus $X_{t}$ remains adequate. Assume that participant $j$ wishes to burn $y$ tokens at time $t>0$. Upon burning, the participant should receive $x=y\cdot r_{t}$ ETH. The transaction is allowed if
$y\leq Y^{(j)}_{t}\text{, }x\leq B_{t},\text{ and }x<X_{t}-SCR_{t}.$ (4)
Provided that (4) is satisfied then we update the contract balance and the surplus as
$B_{t+h}=B_{t}-x,\text{ and }X_{t+h}=X_{t}-x.$
We further update the token balance of investor $j$ with
@@ 3.3 Withdrawing from the smart contract and burning tokens
3.4 Contract resolution and claim settlement
At time $t=T_{i}$, contract $i$, for some $i=1,\ldots,N_{t}$, is settled as it becomes known whether the event $A_{T_{i}}$ has occurred.
Let us first consider the simple case where $A_{T_{i}}$ does not occur. In this scenario, the number of contracts, liability, and premiums collected are updated as follows,
$N_{t+h}=N_{t}-1,\quad L_{t+h}=L_{t}-l_{i},\quad\Pi_{t+h}=\Pi_{t}-(1+\eta_{S_{i}})\cdot l_{i}\cdot\theta_{i}.$
Recall that $S_{i}$ is the time at which contract $i$ was underwritten. The SCR and MCR are also updated as
$\text{SCR}_{t+h}=\text{Quantile}(L_{t+h}-\Pi_{t+h};\alpha_{\text{SCR}}),\quad\text{MCR}_{t+h}=\text{Quantile}(L_{t+h}-\Pi_{t+h};\alpha_{\text{MCR}}),$
using the recursive method based on a normal approximation of the distribution of $L_{t}$ described in Section 4. The balance of the contract remains unchanged $B_{t+h}=B_{t}$, while the surplus is updated as
$X_{t+h}=X_{t}+(1+\eta_{S_{i}})\cdot\theta_{i}\cdot l_{i}.$
The premium for contract $i$ is now considered "earned." The exchange rate is updated as
$r_{t+h}=\frac{X_{t+h}}{\bar{Y}_{t+h}},$
@@ 3.4 Contract resolution and claim settlement
If $A_{T_{i}}$ occurred, the balance, number of contracts, liability, and premiums collected are updated as follows,
$\widetilde{N}_{t}=N_{t}-1,\quad\widetilde{L}_{t}=L_{t}-l_{i},\quad\widetilde{\Pi}_{t}=\Pi_{t}-(1+\eta_{S_{i}})\cdot l_{i}\cdot\theta_{i}.$
The SCR and MCR are also updated as
$\widetilde{\text{SCR}}_{t}=\text{Quantile}(\widetilde{L}_{t}-\widetilde{\Pi}_{t};\alpha_{\text{SCR}}),\quad\widetilde{\text{MCR}}_{t}=\text{Quantile}(\widetilde{L}_{t}-\widetilde{\Pi}_{t};\alpha_{\text{MCR}}).$
of the contract and the surplus are updated as
$\widetilde{B}_{t}=B_{t}-l_{i},\text{ and }\widetilde{X}_{t}=X_{t}-l_{i}+(1+\eta_{S_{i}})\cdot\theta_{i}\cdot l_{i}.$
The use of the tilde indicates that this update is not necessarily final, as the surplus may cross the MCR threshold
$\widetilde{X}_{t}\leq\text{MCR}_{t+h}.$
If this occurs, the smart contract is reset. The outstanding balance is first used to reimburse the policyholders of ongoing contracts. The premium refunded to the policyholder is the amount paid, excluding the participation loading that granted tokens to the policyholder. The total amounts is therefore given by
$x=\sum_{i=1}^{\widetilde{N}_{t}}\mathbb{I}_{\xi^{(i)}_{t}=0}\left(1+\eta_{S_{i}}^{(1)}\right)\cdot\theta_{i}\cdot l_{i}.$
If $\widetilde{X}_{t}>x$, the premiums are fully refunded, and the remaining wealth, $\widetilde{X}_{t}-x$, is distributed to the investors according to their token holdings, using the updated exchange rate
@@
The smart contract transfers
@@
to all investors $j=1,\ldots,M_{t}$.
If $\widetilde{X}_{t}\leq x$, the policyholders are reimbursed pro rata based on the premiums they paid. Specifically, we define weights for the policyholders with active contracts as
$w_{i}=\frac{\mathbb{I}_{\xi^{(i)}_{t}=0}\left(1+\eta_{S_{i}}^{(1)}\right)\cdot\theta_{i}\cdot l_{i}}{\sum_{i=1}^{\widetilde{N}_{t}}\mathbb{I}_{\xi^{(i)}_{t}=0}\left(1+\eta_{S_{i}}^{(1)}\right)\cdot\theta_{i}\cdot l_{i}},$
and transfer
@@
to policyholder $i$ for $i=1,\ldots\widetilde{N}_{t}$.
After completing these transfers, some smart contract variables are reset to their initial state with
$B_{t+h}=0,\quad X_{t+h}=0,\quad L_{t+h}=0,\quad\Pi_{t+h}=0,\quad\text{SCR}_{t+h}=0\quad\text{MCR}_{t+h}=0\quad r_{t+h}=1.$
@@
$Y^{(j)}_{t+h}=0,\quad\text{for all }j=1,\ldots,M_{t+h}.$
@@
The practical implementation of the smart contract using the programming language solidity is documented in the online accompaniement of this paper [^3] https://github.com/LaGauffre/smart_parametric_insurance/blob/main/latex/sup_material_blockchain_parametric_insurance.pdf.
4 Recursive calculation of the solvency capitals
Recall that the liability at time $t>0$ is given by
$L_{0}=0,\quad L_{t}=\sum_{i=1}^{N_{t}}Z_{i}\mathbb{I}_{S_{i}\leq t,T_{i}>t},\quad\text{for }t\geq 0,$ (5)
where $Z_{i}=l_{i}\mathbb{I}_{A_{T_{i}}}$ for $i\geq 0$. Assume that the Bernouilli variables $\left(\mathbb{I}_{A_{T_{i}}}\right)_{i\geq 0}$ are mutually independent. Then, provided that $\sum_{i=1}^{N_{t}}\mathbb{I}_{S_{i}\leq t,T_{i}>t}=n$, the distribution of $L_{t}$ corresponds to that of the aggregate losses in the individual risk model described in standard actuarial textbooks such as Klugman [18, Chapter 9, Section 11]. While the exact distribution may be computed via combinatorial analysis, this approach becomes impractical as $n$ grows large. A more efficient method is to use a recursive algorithm like that of De Pril [8]. 1 recalls this algorithm adapted to our notations.
Proposition 1.
Assume that $l_{i}\in\mathbb{N}$ for $i=1,\ldots,n$ then we have
$\mathbb{P}(L_{t}=0)=\prod_{i=1}^{n}(1-p_{i}),$
and
$\mathbb{P}(L_{t}=x)=\sum_{i=1}^{n}v_{i}(x),$
where
$v_{i}(x)=\begin{cases}0,&x<l_{i},\\ \frac{p_{i}}{1-p_{i}}l_{i}\left[\mathbb{P}(L_{t}=x-l)-v_{i}(x-l_{i})\right],&\text{ otherwise,}\end{cases}$
for $i=1,\ldots,n$.
Once the probability distribution of $L_{t}$ is obtained, we can compute its cumulative distribution function and use a bisection algorithm to determine the quantile of interest. While this method is significantly more efficient than combinatorial analysis, it can still be time-consuming for large $n$ and large values of $l_{i}$. Moreover, numerical instabilities are common in such recursive formulas, often arising from operations on very small probabilities or the subtraction of nearly equal values. In such cases, one can consider approximation methods, such as the approach proposed by Hipp [17], which uses a compound Poisson distribution as an approximation. The distribution of the compouund Poisson distribution then requires the use of Panjer’s algorithm of Panjer [23] or a fast Fourier transform inversion method, see Embrechts and Frei [11] for a comparison with Panjer’s recursion.
The numerical procedures mentioned above are effective but not very practical for automated implementation on the blockchain. As noted in the introduction, smart contract functions cannot encode complex operations. This limitation means that solvency capital calculations would need to be performed off-chain, requiring intervention from the contract manager. Moreover, any change in the portfolio at time $s\geq t$ would necessitate recomputation of the probability distribution of the liability $L_{s}$ to determine $\text{SCR}_{s}$ or $\text{MCR}_{s}$.
For instance, if a policy $(A_{T},l)$ is added or removed (e.g., due to settlement, in which case $s=T$), we would need a method to update the SCR efficiently. To achieve this, we rely on a normal approximation, justified by a generalized central limit theorem, which applies to independent random variables with finite variance. Denote by
$\mathbb{E}(Z_{i})=\mu_{i}=\theta_{i}\cdot l_{i},\text{ and }\mathbb{V}(Z_{i})=\sigma_{i}^{2}=\theta_{i}\cdot(1-\theta_{i})\cdot l_{i}^{2}\text{ for }i=1,\ldots,n,$
the mean and variance of the $Z_{i}$’s. Let $s_{n}^{2}=\sum_{i=1}^{n}\sigma_{i}^{2}$ be the sum of the variances. Suppose that the sequence of random variables $(Z_{i})_{i=1,\ldots,n}$ satisfies
$\underset{n\rightarrow\infty}{\lim}\frac{1}{s_{n}^{2}}\sum_{i=1}^{n}\mathbb{E}\left[\left(Z_{i}-\mu_{i}\right)^{2}\mathbb{I}_{|Z_{i}-\mu_{i}|>\epsilon s_{n}^{2}}\right]=0,\text{ for every }\epsilon>0,$ (6)
then we have
$\frac{1}{s_{n}}\sum_{i=1}^{n}(Z_{i}-\mu_{i})\overset{\mathcal{D}}{\longrightarrow}\mathsf{Normal}(0,1),\text{ as }n\rightarrow\infty,$
where $\mathsf{Normal}(0,1)$ is the standard normal distribution and $\overset{\mathcal{D}}{\longrightarrow}$ means convergence in distribution.
Proposition 2.
Suppose that the $l_{i}$’s are bounded, then condition 6 holds.
Proof.
The Lyapunov condition holds if there exists $\delta>0$ such that
$\underset{n\rightarrow\infty}{\lim}\frac{1}{s_{n}^{2+\delta}}\sum_{i=1}^{n}\mathbb{E}\left(|Z_{i}-\mu_{i}|^{2+\delta}\right)=0,$ (7)
and implies condition (6). Denote by $l=\sup_{i\geq 0}l_{i}$ then we have
$\frac{1}{s_{n}^{2+\delta}}\sum_{i=1}^{n}\mathbb{E}\left(|\xi_{i}-\mu_{i}|^{2+\delta}\right)$$\leq$ $\frac{l^{\delta}}{s_{n}^{2+\delta}}\sum_{i=1}^{n}\mathbb{E}\left(|\xi_{i}-\mu_{i}|^{2}\right)$
$=$ $\left(\frac{l}{s_{n}}\right)^{\delta}\rightarrow 0\text{ for every }\delta>0.$
We note that
$L_{t}-\Pi_{t}=\sum_{i}^{n}Z_{i}-\sum_{i=1}^{n}(1+\eta_{S_{i}})\cdot\mu_{i}=s_{n}\left(\frac{1}{s_{n}}\sum_{i}^{n}(Z_{i}-\mu_{i})\right)-\sum_{i=1}^{n}\eta_{S_{i}}\cdot\mu_{i}.$
We deduce the following approximations of the SCR and MCR,
$\text{SCR}_{t}\approx s_{n}\cdot q_{\alpha_{\text{SCR}}}-\eta_{S_{i}}\cdot\sum_{i=1}^{n}\mu_{i},\text{ and }\text{MCR}_{t}\approx s_{n}\cdot q_{\alpha_{\text{MCR}}}-\eta_{S_{i}}\cdot\sum_{i=1}^{n}\mu_{i},$
where $q_{\alpha}$ is the quantile of order $\alpha$ of the standard normal distribution. Suppose that at time $t=S$, a contract $(A_{T},l)$ such that $\mathbb{P}(A_{T})=\theta$, is underwritten then we update the SCR as
$\text{SCR}_{t+h}=\sqrt{s_{n}^{2}+\theta\cdot(1-\theta)\cdot l^{2}}\cdot q_{\alpha_{\text{SCR}}}-\left(\sum_{i=1}^{n}\eta_{S_{i}}\mu_{i}+\eta_{S}\cdot\theta\cdot l\right).$
Conversely, suppose that at time $t=T_{k}$ contract $k\in\{1,\ldots,N_{t}\}$ is settled then we update the SCR as
$\text{SCR}_{t+h}=\sqrt{s_{n}^{2}-\theta_{k}\cdot(1-\theta_{k})\cdot l_{k}^{2}}\cdot q_{\alpha_{\text{SCR}}}-\left(\sum_{i=1}^{n}\eta_{S_{i}}\mu_{i}-\eta_{S_{k}}\cdot\theta_{k}\cdot l_{k}\right).$
Note that updating the MCR is performed in the same manner, with $q_{\alpha_{\text{SCR}}}$ replaced by $q_{\alpha_{\text{MCR}}}$. The updates to the SCR and MCR can be implemented recursively in Solidity, as they only require updating $s_{n}^{2}$ and $\sum_{i}\eta_{S_{i}}\mu_{i}$. To achieve this, we implemented a custom square root function in Solidity; details can be found in the online supplementary materials [^4] https://github.com/LaGauffre/smart_parametric_insurance/.
The quantiles of the normal distribution used to determine the SCR and MCR are hardcoded but can be modified by the contract owner upon recommendation from the board of token holders.
Remark 4.1.
We have not addressed the estimation of $\theta$ in detail. Standard actuarial practice would recommend using historical data and a regression approach to account for the characteristics of the insured event. Logistic regression is particularly suitable for this task, especially if the probability of the event’s occurrence is influenced by covariates. This is often the case for weather-related events, as discussed in Figueiredo et al. [13]. The formula for calculating $\theta$ could be hardcoded into the smart contract, with covariate values retrieved via relevant APIs, commonly referred to as data oracles in the blockchain ecosystem. Additionally, a function could be included to allow the contract owner to update the regression coefficients if needed.
In cases where the smart contract only underwrites the same type of risk—meaning the Bernoulli variables are independent and identically distributed—a simple beta-binomial credibility model can be used for experience rating; see the monograph of Hans Bühlmann [16, Section 2.1.2] for details.
In our implementation, however, we do not query any API nor include a pricing formula within the smart contract. Instead, we emulate this by generating a random number between 00 and $1$ based on an event description represented as a string. We apply the keccak256 hash function to the string to produce a 256-bit number, which is then converted into a binary sequence of length 256. The resulting number is taken modulo $10{,}000$, and the outcome is divided by $10{,}000$, yielding a pseudo-random probability drawn from a standard uniform distribution.
To determine whether compensation is paid, we process the description of the observed event in the same manner. If the resulting number for the observed event is smaller than that of the original event, a compensation payment is triggered. The getRandomNumber function is provided in the online accompaniement.
A brief illustration of the performance of our normal approximation for estimating the MCR and SCR is provided in 1.
Example 1.
Assume that a collection of contracts $\{(A_{T_{i}},l_{i})\}_{i=1,\ldots,n}$ are underwritten sequentially at times $S_{1}\leq S_{2}\leq\ldots\leq S_{n}$ such that $S_{n}\leq\underset{i=1,\ldots,n}{\min}T_{i}$. Suppose that the probability of the events are given by
$\mathbb{P}(A_{T_{i}})=\theta_{i}\sim\mathsf{Unif}(0,1)\text{ for }i=1,\ldots,n.$
to be consistent with the probability generating process described in Remark 4.1. We also randomly generate the potential compensations with
$l_{i}\sim\mathsf{Unif}(\{1,5\})\text{ for }i=1,\ldots,n,$

### v2

@@ 4.3 Method of the smart contract
The state of a smart contract changes according to the actions of blockchain users. In our parametric insurance example, these users include the contract owner, surplus providers, and policyholders (see Section 4.1). Participants interact with the smart contract through specific functions, known as methods, which modify the system’s state. The remainder of this section describes each possible action and its impact on the stochastic processes introduced thus far.
4.3.1 Funding the contract
@@
Assume that an investor transfers $x$ ETH to the contract at time $t>0$. As a result, both the balance and the surplus of the contract increase as follows:
@@
where $h>0$ represents an infinitesimal time lapse.
If the investor is new, the count of distinct addresses increases, such that $A_{t+h}=A_{t}+1$; otherwise, $A_{t+h}=A_{t}$. The $j$-th component of the vector $(Y_{t})_{t\geq 0}$, where $j\in\{1,\ldots,A_{t+h}\}$, is updated as:
@@
4.3.2 Underwriting a parametric insurance policy
A customer wishes to purchase an insurance contract $(S,T,Q,\overline{Q},l)$ at time $t=S$ such that $S<S^{\ast}$. The premium for such a parametric insurance contract is given by:
@@ 4.3.2 Underwriting a parametric insurance policy
where $\theta=\mathbb{P}(Q_{T}>\overline{Q}|\mathcal{F}_{S})$. The contract can only be underwritten if the surplus of the smart contract is sufficient to cover the resulting increase in the Solvency Capital Requirement (SCR).
The liability, the cumulative premium collected, and the number of active model points would then be updated to:
$\widetilde{\Lambda}_{t}=\Lambda_{t}+l,\quad\widetilde{L}_{t}=L_{t}+l\cdot\mathbb{I}_{Q_{T}>\bar{Q}},\quad\widetilde{\Pi}_{t+h}=\Pi_{t}+\pi,$
and
$\widetilde{M}^{\text{active}}_{t}=\begin{cases}M^{\text{active}}_{t},&\text{ifthe new contract is grouped into an active model point},\\ M^{\text{active}}_{t}+1,&\text{otherwise}.\end{cases}$
@@
$\widetilde{\text{SCR}}_{t}=\begin{cases}\text{Quantile}(\widetilde{L}_{t}-\widetilde{\Pi}_{t};\alpha_{\text{SCR}}),&\text{ if }\widetilde{M}^{\text{active}}_{t}\geq m,\\ \Lambda_{t},&\text{otherwise.}\end{cases}$
@@
If condition (7) holds, the number of contracts, liability, premium collected, and number of active model points are updated as follows:
$N_{t+h}=N_{t}+1,\quad\Lambda_{t+h}=\Lambda_{t}+l,\quad L_{t+h}=L_{t}+l\cdot\mathbb{I}_{Q_{T}>\bar{Q}},\quad\Pi_{t+h}=\Pi_{t}+(1+\eta_{t})\cdot l\cdot\theta\text{, and }M_{t+h}^{\text{active}}=\widetilde{M}^{\text{active}}_{t}$
@@
$\text{SCR}_{t+h}=\begin{cases}\text{Quantile}(L_{t+h}-\Pi_{t+h};\alpha_{\text{SCR}}),&\text{ if }M^{\text{active}}_{t+h}\geq m,\\ \Lambda_{t},&\text{otherwise,}\end{cases}$
and
$\text{MCR}_{t+h}=\begin{cases}\text{Quantile}(L_{t+h}-\Pi_{t+h};\alpha_{\text{MCR}}),&\text{ if }M^{\text{active}}_{t+h}\geq m,\\ \Lambda_{t},&\text{otherwise.}\end{cases}$
@@
to the vector $\Xi_{t}$ that gathers the statuses of the insurance policies. The balance of the contract is updated as:
$B_{t+h}=B_{t}+(1+\eta_{t})\cdot\theta\cdot l.$
4.3.3 Withdrawing from the smart contract and burning tokens
An investor $j\in\{1,\ldots,A_{t}\}$ may withdraw some of her funds by burning protocol tokens at time $t>0$. This action is subject to the conditions that the participant holds enough protocol tokens $Y_{t}^{(j)}$, the smart contract has a sufficient balance $B_{t}$, and the surplus $X_{t}$ remains adequate.
Assume that participant $j$ wishes to burn $y$ tokens at time $t>0$. Upon burning, the participant should receive $x=y\cdot r_{t}$ ETH. The transaction is allowed if the following conditions are met:
$y\leq Y_{t}^{(j)},\quad x\leq B_{t},\quad\text{and}\quad x<X_{t}-\text{SCR}_{t}.$ (8)
Provided that the conditions in (8) are satisfied, we update the contract balance and the surplus as follows:
$B_{t+h}=B_{t}-x,\quad\text{and}\quad X_{t+h}=X_{t}-x.$
We further update the token balance of investor $j$ with:
@@ 4.3.3 Withdrawing from the smart contract and burning tokens
Remark 4.1.
Since the events that trigger compensation are predictable in the short term, it is essential to impose timing constraints on investors’ fund withdrawals. One practical solution would be to require investors to submit withdrawal orders, which the smart contract owner would validate before processing. Under this approach, investors would indicate their intention to withdraw funds at a future date and specify the desired amount in advance. Note that this feature has not been implemented in our current project.
4.3.4 Contract resolution and claim settlement
At time $t=T_{i}$, the contract $(S_{i},T_{i},Q_{i},\overline{Q}_{i},l_{i})$, for some $i=1,\ldots,N_{t}$, is settled as it becomes known whether the event $Q_{T_{i}}>\overline{Q}_{i}$ has occurred. The liabilities and the premum collected are updated as follows:
$\widetilde{\Lambda}_{t}=\Lambda_{t}-l_{i}\text{, }\widetilde{L}_{t}=L_{t}-l_{i},\quad\widetilde{\Pi}_{t}=\Pi_{t}-(1+\eta_{S_{i}})\cdot l_{i}\cdot\theta_{i}.$
The balance of the contract remains unchanged, $\widetilde{B}_{t}=B_{t}$, while the surplus is updated as:
$\widetilde{X}_{t}=\begin{cases}X_{t}+(1+\eta_{S_{i}})\cdot\theta_{i}\cdot l_{i},&\text{ if }Q_{T_{i}}\leq\overline{Q}_{i},\\ X_{t}+(1+\eta_{S_{i}})\cdot\theta_{i}\cdot l_{i}-l_{i},&\text{ otherwise}.\end{cases}$
Recall that $S_{i}$ is the time at which contract $i$ was underwritten. The number of active model poinst is updated as follows:
$\widetilde{M}^{\text{active}}_{t}=\begin{cases}M^{\text{active}}_{t}-1,&\text{if the contract was the last contract of a given model point},\\ M^{\text{active}}_{t},&\text{otherwise}.\end{cases}$
The Solvency Capital Requirement (SCR) and Minimum Capital Requirement (MCR) are also updated as:
$\widetilde{\text{SCR}}_{t}=\begin{cases}\text{Quantile}(\widetilde{L}_{t}-\widetilde{\Pi}_{t};\alpha_{\text{SCR}}),&\text{ if }\widetilde{M}^{\text{active}}_{t}\geq m,\\ \widetilde{\Lambda}_{t},&\text{otherwise,}\end{cases}$
and
$\widetilde{\text{MCR}}_{t}=\begin{cases}\text{Quantile}(\widetilde{L}_{t}-\widetilde{\Pi}_{t};\alpha_{\text{MCR}}),&\text{ if }\widetilde{M}^{\text{active}}_{t}\geq m,\\ \widetilde{\Lambda}_{t},&\text{otherwise.}\end{cases}$
At this stage we distinguish two cases. First assume that $\widetilde{X}_{t}>\widetilde{\text{MCR}}_{t}$ then all the processes are updated for good with
$B_{t+h}=\widetilde{B}_{t}\text{, }X_{t+x}=\widetilde{X}_{t}\text{, }\Lambda_{t+h}=\widetilde{\Lambda}_{t}\text{, }L_{t+h}=\widetilde{L}_{t}\text{, }\Pi_{t+x}=\widetilde{\Pi}_{t}\text{, }\text{SCR}_{t+h}=\widetilde{SCR}_{t}\text{, }\text{MCR}_{t+x}=\widetilde{\text{MCR}}_{t}\text{, and }M_{t+h}^{active}=\widetilde{M}^{\text{active}}_{t}.$
The premium for contract $i$ is now considered "earned." The exchange rate is updated as:
$r_{t+h}=\frac{X_{t+h}}{\overline{Y}_{t+h}},$
@@ 4.3.4 Contract resolution and claim settlement
Now suppose that the surplus downcrosses the MCR as $\widetilde{X}_{t}\leq\widetilde{\text{MCR}}_{t}$. If this occurs, the smart contract is reset. The outstanding balance is first used to reimburse the policyholders of active contracts whose status becomes "cancelled". The total amount is therefore given by:
$x=\sum_{i=1}^{N_{t}}\mathbb{I}_{\xi^{(i)}_{t}=0}\left(1+\eta_{S_{i}}\right)\cdot\theta_{i}\cdot l_{i}.$
If $\widetilde{X}_{t}>x$, the premiums are fully refunded, and the remaining wealth, $\widetilde{X}_{t}-x$, is distributed to the investors according to their token holdings, using the updated exchange rate:
@@
The smart contract transfers:
@@
to all investors $j=1,\ldots,A_{t}$.
If $\widetilde{X}_{t}\leq x$, the policyholders are reimbursed pro rata based on the premiums they paid. Specifically, we define weights for the policyholders with active contracts as:
$w_{i}=\frac{\mathbb{I}_{\xi^{(i)}_{t}=0}\left(1+\eta_{S_{i}}\right)\cdot\theta_{i}\cdot l_{i}}{\sum_{i=1}^{N_{t}}\mathbb{I}_{\xi^{(i)}_{t}=0}\left(1+\eta_{S_{i}}\right)\cdot\theta_{i}\cdot l_{i}},$
and transfer:
@@
to the policyholder associated to policy $i$ for $i=1,\ldots,N_{t}$.
After completing these transfers, the smart contract variables are reset to their initial state with:
$B_{t+h}=0,\quad X_{t+h}=0,\quad\lambda_{t+h}=0,\quad L_{t+h}=0,\quad\Pi_{t+h}=0,\quad\text{SCR}_{t+h}=0,\quad\text{MCR}_{t+h}=0,\quad r_{t+h}=1\text{, and }M_{t+h}^{active}=0.$
@@
$Y^{(j)}_{t+h}=0,\quad\text{for all }j=1,\ldots,A_{t+h}.$
@@
The practical implementation of the smart contract using the programming language Solidity is documented in the online accompaniment of this paper [^3] https://github.com/LaGauffre/smart_parametric_insurance/blob/main/latex/sup_material_blockchain_parametric_insurance.pdf.
5 Interaction with the smart contract
During the development phase, a smart contract is not immediately deployed on the Ethereum mainnet. Instead, developers use testnets to simulate the behavior of the real blockchain, ensuring that the smart contract operates as intended. For our project, we deployed the smart contract on the Ethereum testnet known as Sepolia. Sepolia includes a block explorer called Etherscan [^4] See https://sepolia.etherscan.io/, which provides an API for data retrieval. Our contract, named InsuranceLogic, is located at the address: